What is the slope and y-intercept of the line, of the equations y= -9/7x + 2 and 9x+3y=6?
Answer by rfer(16322)

A)
slope m=-9/7
y int=2 (0,2)
B)
9x+3y=6
3y=-9x+6
y=-3x+2
slope m=-3
y int=2 (0,2)